OpenStudy (anonymous):

in the 2004 summer olympic games in athens,greece, the united states athletes won 2 more than 3 times the number of gold metals won by the french athletes. the united won 24 more god metals than the french. solve the equation 24+f=3f+2 to find the number of gold metals won by the french athletes

OpenStudy (anonymous):

3f-f = 24-2 f =22/2 =11